In a random mating population of 20000, 16800 individuals are the dominant then how many individuals will be homozygous?
1
Expert's answer
2018-08-14T05:53:48-0400
The number of individuals with recessive phenotype is 20000 - 16800 = 3200.
Their proportion in the population is 3200 / 20000 = 0.16
As individuals with recessive phenotype can have aa genotype only, the frequency of recessive allele is sqrt (0.16)=0.4
The frequency of dominant allele is 1 - 0.4=0.6
The frequency of homozygous individuals with dominant phenotype is 0.6^2 = 0.36
The number of such individuals is 0.36 * 20000 = 7200